as2s3+nano3+na2co3 -> na3aso4 + na2so4 + nano2 +co2 determine the theoretical mass of reactants needed to form 50 g of sodium arsenate
1
Expert's answer
2020-12-01T14:23:47-0500

As2S3+14NaNO3+6Na2CO3 -> 2Na3AsO4 + 3Na2SO4 + 14NaNO2 +6CO2

M(Na3AsO4)=208 g/mol

n (Na3AsO4)=m/M = 50/208 = 0.24 mol

M (As2S3) = 246 g/mol

M (NaNO3) = 85 g/mol

M (Na2CO3) = 106 g/mol


n (As2S3) = 1/2 x n (Na3AsO4) = 1/2 x 0.24 = 0.12 mol

m (As2S) = n x M = 0.12 x 246 = 29.52 g


n (NaNO3) = 14/2 x n (Na3AsO4) = 14/2 x 0.24 = 1.68 mol

m (NaNO3) = n x M = 1.68 x 85 = 142.8 g


n (Na2CO3) = 6/2 x n (Na3AsO4) = 6/2 x 0.24 = 0.72 mol

m (Na2CO3) = n x M = 0.72 x 106 = 76.32 g